OpenStudy (anonymous):

solve. 3/6= k/18 = k

OpenStudy (anonymous):

@Bee_see

OpenStudy (anonymous):

cross multiply

OpenStudy (sleepyjess):

These 2 expressions are equivalent. What has been done to 6 to get to 18?

OpenStudy (anonymous):

3/6 = 9/18

OpenStudy (sleepyjess):

@mathkid911 , Please do not give out direct answers.

OpenStudy (anonymous):

K=9

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Because you cross multiply

OpenStudy (bee_see):

9. 3 x 18=54. 54/6=9

OpenStudy (anonymous):

Yet now I get it thanks ^-^

OpenStudy (sleepyjess):

@PurpleGirl138 , Please do not give out direct answers.

OpenStudy (anonymous):

I have 3 more

OpenStudy (sleepyjess):

Ok, open a new question and I would be happy to help :)

OpenStudy (anonymous):

@sleepyjess I didnt, I told him to cross multiply lol